OpenStudy (aravindg):

from each corner of a square of tin ,30cm on a side ,,small squares of side x cm are removed and the edges turned up to form an open box..Express the volume V (m^3) as a function of x and examine the range of each variable..........

OpenStudy (aravindg):

answer plzzzzzzzzzzzzzzz

OpenStudy (anonymous):

here side of square = 30 cms now after removing x cms from the corners we have base = 30-2x and, height =x so, volume v= (30-2x)^2x cms^3 so, volume v= (30-2x)^2x /1000 m^3